### v1

# MMRel: A Relation Understanding Dataset and Benchmark in the MLLM Era

Abstract
Despite the recent advancements in Multi-modal Large Language Models (MLLMs), understanding inter-object relations, i.e., interactions or associations between distinct objects, remains a major challenge for such models. This issue significantly hinders their advanced reasoning capabilities and is primarily due to the lack of large-scale, high-quality, and diverse multi-modal data essential for training and evaluating MLLMs. In this paper, we provide a taxonomy of inter-object relations and introduce Multi-Modal Relation Understanding (MMRel), a comprehensive dataset designed to bridge this gap by providing large-scale, high-quality and diverse data for studying inter-object relations with MLLMs. MMRel features three distinctive attributes: (i) It includes over 15K question-answer pairs, which are sourced from three distinct domains, ensuring large scale and high diversity; (ii) It contains a subset featuring highly unusual relations, on which MLLMs often fail due to hallucinations, thus are very challenging; (iii) It provides manually verified high-quality labels for inter-object relations. Thanks to these features, MMRel is ideal for evaluating MLLMs on relation understanding, as well as being used to fine-tune MLLMs to enhance relation understanding and even benefit overall performance in various vision-language tasks. Extensive experiments on various popular MLLMs validate the effectiveness of MMRel. Both MMRel dataset and the complete labeling scripts have been made publicly available.
1 Introduction
Multi-Modal Large Language Models (MLLMs) [1, 2, 3, 4, 5, 6, 7] have demonstrated impressive capabilities in various vision-language tasks [8, 9, 10, 11]. One key step in these tasks is to comprehend inter-object relations, which are defined by the interactions or associations between distinct objects in images [12]. Understanding these relations is crucial for MLLMs to perform advanced reasoning, provide accurate responses in visual question answering [8, 9], and generate detailed image captions [10, 11]. However, recent studies [13, 14, 15, 16] have shown that most existing MLLMs struggle with handling inter-object relations, as illustrated in Fig. 1. The primary reasons for this challenge include the lack of a formal definition of inter-object relations and the absence of benchmarks that feature large-scale, diverse, and high-quality vision-language data.
To fill up the gap by building a comprehensive dataset for MLLMs’ relation understanding, we first partition inter-object relations into three distinct categories, namely, spatial relations (e.g., dog is left to cat), action relations (e.g., boy eats pizza), and comparative relations (e.g., apple is smaller than watermelon). Furthermore, we define the term relation hallucinations by scenarios where the MLLM-generated inter-object relations align with common sense but deviate from the factual contents in images. For example, MLLMs responds that “a man drives a car” while the ground truth is “a man pushes a car”. In addition, we examine pioneer benchmarks that were not intended for but did allow assessing MLLMs’ relation understanding capabilities [17, 18, 19, 20, 21, 22, 23, 24, 25]. As Tab. 1 shows, most existing benchmarks are typically small-scale with homogeneous data, capture one or two categories of inter-object relations only, and lack diversity with images from a single domain.
Table 1: A comparison of MMRel with existing relation understanding benchmarks. MMRel is featured with large-scale, diverse, and high-quality data on inter-object relations.
Dataset Scale Taxonomy Domain MLLM+++ Adv- Evaluation
Spatial Action Comparative Real SD Dall-E Human subset
MMHAL [18] 12 $\checkmark$ $\times$× $\times$× $\checkmark$ $\times$× $\times$× $\checkmark$ $\times$× LLM
RAH [21] 500 $\checkmark$ $\checkmark$ $\times$× $\checkmark$ $\times$× $\times$× $\times$× $\times$× Yes/No
MERLIM [22] - $\checkmark$ $\times$× $\times$× $\checkmark$ $\times$× $\times$× $\times$× $\times$× Yes/No
FAITHSCORE [23] - - - - $\checkmark$ $\times$× $\times$× $\checkmark$ $\times$× LLM
M-HalDetect [20] - - - - $\checkmark$ $\times$× $\times$× $\checkmark$ $\times$× Human
AMBER [17] 1.7K $\checkmark$ $\times$× $\times$× $\checkmark$ $\times$× $\times$× $\checkmark$ $\times$× Yes/No
SPEC [19] 3.5K $\checkmark$ $\times$× $\checkmark$ $\times$× $\checkmark$ $\times$× $\times$× $\times$× Choice
Hallusion [25] 150 $\times$× $\times$× $\checkmark$ $\checkmark$ $\times$× $\times$× $\times$× $\times$× Yes/No
MME [19] 60 $\checkmark$ $\times$× $\times$× $\checkmark$ $\times$× $\times$× $\times$× $\times$× Yes/No
MMRel 15K $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ Yes/No
Figure 1: MLLMs often fail to understand inter-object relations.
With the definition of inter-object relations and relation hallucinations, we create Multi-Modal Relation Understanding (MMRel), a large-scale benchmark that comprises multi-modal data of three categories of relations (i.e. spatial, action, and comparative) that are sourced from three distinctive domains (i.e., real images, synthetic images from SDXL [26], and synthetic images from Dall-E [27]). We present sample images from MMRel in Fig. 2. To ensure the quality of the collected data, we adopt a semi-automatic pipeline that leverages MLLMs [27, 28] to generate images and annotations based on textual prompts and then verifies and corrects the generated images and annotations by human reviewers. As MLLMs are known to have the issue of hallucination (especially on relation understanding tasks), to evaluate and prevent hallucinations, we create a challenging adversarial subset in MMRel which utilizes relations that deviate from common sense to assess the relation understanding capabilities of MLLMs rigorously. All these lead to a benchmark that comprises over 15,000 question-answer pairs, each of which comes with clearly defined annotations and relation categories. To make the evaluation straightforward and easy to understand, we adopt a Yes/No framework, which has been widely used in previous datasets [17, 21, 22, 24, 25, 29].
Figure 2: Illustration of MMRel samples. MMRel covers three categories of inter-object relations, namely, spatial (orange), action (blue), and comparative (green), and its images are sourced from three domains of different styles: (a) natural images, (b) synthetic images by SDXL [26], and (c) synthetic images by Dall-E [27]. It is also featured by a challenging subset that comprise images with unusual relations as in (c). More Dall-E samples are provided in the Appendix.
MMRel can serve different purposes thanks to its large-scale, diverse, and high-quality vision-language data. One typical scenario is evaluating the relation understanding capabilities of existing MLLMs. We conduct comprehensive evaluations on this aspect, revealing that both existing MLLMs [2, 5, 7] and hallucination mitigation techniques [30, 31, 32] exhibit sub-optimal performance in understanding inter-object relations. Another typical scenario is instruction tuning of MLLMs for understanding inter-object relations [1, 2]. Extensive experiments verify the effectiveness of MMRel as both evaluation set and training data, and also demonstrate that MMRel data is beneficial not only to relation understanding but also to other perceptual tasks. All these studies affirm MMRel’s values in exploring more powerful MLLMs in terms of improving advanced perception.
The contributions of this work can be summarized in three aspects. First, it systematically defines inter-object relations as well as relation hallucinations under the context of MLLMs. This lays a solid foundation for future study of inter-object relations and hallucination mitigation. Second, we create MMRel, a large-scale benchmark that comprises over 15K high-quality vision-language data that are sourced from diverse domains with clearly defined inter-object relations. To the best of our knowledge, this is the first large-scale benchmark on inter-object relations that comes with diverse and high-quality vision-language data. Third, extensive experiments demonstrate the value of MMRel for both evaluating existing MLLMs on relation understanding and fine-tuning for the enhanced power of relation understanding.
2 Reviews on Existing Datasets for Relation Understanding
Research on inter-object relations boasts a rich history, with Scene Graph Generation (SGG) [33, 34] serving as a foundational area of focus. SGG is specifically dedicated to identifying objects and their interrelations within images. Leveraging specialized object detection methods [35, 36] that excel in detecting both tiny and inconspicuous objects, SGG datasets such as Visual Genome (VG) [12] are distinguished by their extremely detailed and comprehensive annotations. Following the recent successes [37], Multi-Modal Large Language Models (MLLMs) have received significant attention. In response, new datasets [14, 38] have been developed to assess the capabilities of MLLMs in relation understanding.
Figure 3: Comparing MMRel with existing datasets. Existing datasets exhibit the following limitations: (a)-(b) negative choices for action relations are implausible, thus can be easily ruled out; (c) evaluation metrics are complex and subjective; (d) definitions and taxonomies of relations are ambiguous. MMRel mitigates the limitations of existing datasets by enhancing the precision and utility of inter-object relation data.
VL-Checklist [13] and ARO [38] are representative datasets that evaluate the capabilities of MLLMs to understand fine-grained information, including inter-object relations. However, these datasets inherit images and labels directly from VG [12] and adopt a binary Positive/Negative evaluation framework, as illustrated in Fig. 3(a)-(b). This evaluation approach has significant limitations, as it generates negative choices randomly without human verification. Specifically, the negative choices are generated by: (i) randomly altering the relations (e.g. man riding shirt); (ii) simply reversing the order of objects (e.g. grass is eating the horse), making such negative options semantically implausible or even impossible. Given this limitation, we hypothesize that MLLMs can easily deduce the correct answers even without seeing images, as the advanced capabilities of LLM enable MLLMs to effectively dismiss negative choices that starkly deviate from common sense.
Moreover, although recent MLLMs [1, 2, 3, 4, 5, 6, 7] are versatile, their proficiency in object detection is limited due to pre-training on coarse granularity image-text pairs [39, 40]. Therefore, the overly dense and occasionally incorrect annotations from existing datasets, such as VG [12], are unsuitable for assessing MLLMs. Consequently, re-labeling some existing dataset is essential for developing new benchmarks that are more suitable for the evaluation of MLLMs’ capabilities in relation understanding. This has triggered the recent initiatives such as M-HalDetect [20] and AMBER [17] that concentrate on generating new labels. Specifically, M-HalDetect [20] incorporates manually annotated relations within lengthy descriptive sentences, with evaluations performed by human assessors (refer to Fig. 3(c)). However, this metric does not straightforwardly reflect the capabilities in relation understanding and introduces potential risks of subjective comparisons. On the other hand, AMBER [17] simplifies the classification of relationships to contact or not and employs a Yes/No evaluation framework (refer to Fig. 3(d)). While this approach is straightforward, it leads to an imprecise evaluation due to the restricted definition of relations.
The limitations of existing benchmarks can thus be summarized: (i) They are small-scale and lack diversity; (ii) Their annotations are of low quality, with negative choices for action relations often being meaningless; (iii) They lack clear definitions and taxonomies of inter-object relations.
3 Multi-Modal Relation Understanding (MMRel) Dataset
As analyzed in Sec. 1 and Sec. 2, existing datasets for relation understanding are generally constrained by their small-scale and homogeneous nature. This motivates us to develop a new benchmark, Multi-Modal Relation Understanding (MMRel), characterized by large-scale, diverse, and high-quality data on inter-object relations. In this section, we first describe the image sources and relation taxonomies in Sec. 3.1. Subsequently, we introduce a Semi-automatic Data Collection pipeline in Sec. 3.2, designed to amass large-scale, diverse, and high-quality inter-object relation data. Finally, we present the statistics and additional details of the MMRel dataset in Sec. 3.3.
3.1 Data Source and Relation Taxonomy
Following previous works [13, 38], we source our real image subset from the Visual Genome (VG) [12] and select synthetic images from the SPEC [19], which is built based on SDXL [26]. Specifically, we use the relative spatial subset from SPEC [19] for spatial relations and the relative size subset for comparative relations. To further diversify the MMRel, we also generate synthetic images via Dall-E [27]. Given the good clarity of annotations in SPEC [19], we adapt their labels to enhance our dataset’s utility for relation understanding. Details regarding annotation and image generation across other domains are outlined in Sec. 3.2.
Moreover, we employ distinct annotation strategies for each category of relations within our taxonomy. For spatial relations, we limit the categories to left, right, on/over, and under, as other descriptors such as around or next to do not accurately capture relative inter-object positions. For action relations, our annotations are based on the capabilities of Multi-Modal Large Language Models (MLLMs), as detailed in Sec. 3.2. Comparative relations are annotated manually.
3.2 Semi-Automatic Data Collection
In this section, we introduce our Semi-automatic Data Collection pipeline (SemiDC), as illustrated in Fig. 4. This pipeline is capable of annotating large-scale existing images and generating a substantial amount of high-quality synthetic images. As discussed in Sec. 2, re-labeling existing images is essential since their original labels are incompatible with MLLMs. To this end, we design SemiDC to generate high-quality relation annotations via GPT-4V [28] (refer to Fig. 4(a)) for large-scale VG dataset [12]. This process is divided into three stages: (i) Pre-processing: We selectively exclude images featuring complex scenes that pose challenges for GPT-4V [28] in generating accurate annotations. Specifically, we utilize the object labels and corresponding bounding boxes from the VG dataset [12] to identify and exclude images that contain more than 10 tiny objects, each occupying less than 1/20 of the total image area; (ii) Re-labeling via GPT-4V [28]: We employ the in-context learning paradigm [4, 41] to use GPT-4V [28] to generate relation annotations. For example, we provide an image and its accurate relation annotations as a reference, together with a query image accompanied by the textual prompt “Please generate the action relationships between salient objects in this image”. GPT-4V [28] will then produce annotations for the query image based on the text prompt. More details can be found in the Appendix; (iii) Human verification: We manually assess and correct the annotations that are generated by GPT-4V [28], to ensure the quality of the collected inter-object relation data.
Figure 4: Pipeline of Semi-automatic Data Collection (SemiDC). (a) Re-labeling images from Visual Genome [12] datasets. (b) Generating synthetic images based on textual prompts.
As MLLMs are known to have a bias towards positive responses [29, 42], we design MMRel to maintain a balance between “Yes” and “No” answers to mitigate potential biases in evaluation metrics [29]. Specifically, negative for spatial and comparative relations can be simply derived by reversing the order of two objects. For instance, “dog is left to cat” is the inverse of “cat is left to dog”. Nevertheless, as discussed in Sec. 2, randomly generated negative choices for action relations are implausible, which can be easily ruled out even without referring to the images. To address this issue, we first adopt GPT-3.5 [43] to generate plausible negative relations that are well aligned with common sense for given objects, and then verify the generated negative relations by manual checking. For instance, a suitable negative relation for “man eats cake” might be generated as “man bakes cake”. This design is a significant improvement over prior works [13, 38], in terms of both training and evaluation.
Furthermore, we introduce generative models for generating more diverse inter-object relation data. The SemiDC pipeline for generating images (refer to Fig. 2(b)) is also divided into three stages: (i) Preparing textual prompts: We prepare textual prompts describing inter-object relations; (ii) Image generation via Dall-E-3 [27]: We utilize the prepared text prompts as inputs for DALL-E-3 [27] to generate corresponding images. For instance, if we want Dall-E [27] to generate an image depicting a crocodile on the dessert, then we should employ the prompt “a photo-realistic picture depicting an object-object relation: crocodile on dessert” as input. These textual prompts define our targeted relations and the visual style of the generated images. We consider four styles, photo-realistic, watercolor, abstract, and oil painting (refer to Fig. 2(c)), to enrich our dataset; (iii) Human verification of generated images: If the generated images do not match the input relations, we will modify the relations instead of re-generating new images. The same as annotating negative choices for VG [12], we utilize GPT-3.5 [43] to generate negative action relations. More details of SemiDC for both relabeling and generation of inter-object relation data are described in the Appendix.
3.3 Statistics of MMRel
Table 2: Statistics of each subset in MMRel.
Domain Real SDXL Dall-E Total
Taxonomy Spatial Action Comparative Total Spatial Comparative Total Spatial Action Total
[…]
Table 3: Number of objects and categories of relations in MMRel.
Object Spatial Action Comparative
855 5 378 9
Tab. 2 shows the statistics of inter-object relation data in MMRel. Specifically, MMRel comprises over 15,000 question-answer pairs across 7 subsets, spanning 3 domains and 3 categories of relations. Concretely, approximately 68% of the data consists of real images, while the remaining 32% comprises synthetic images (i.e., SDXL [26] and Dall-E [27]). In terms of relation categories, MMRel includes a larger scale of spatial and action relations compared to comparative relations. Thanks to the open-vocabulary capability of GPT-4V [28], MMRel guarantees a diverse range of objects and action relations, as shown in Tab. 3. Regarding comparative relations, we manually re-label VG images, resulting in a subset that, though relatively small in scale with limited diversity, ensures accurate annotations. It is important to note that our annotations strictly describe the objects and their relations, without including any attributes, so that the evaluation can accurately reflect the relation understanding capabilities. For example, if we introduce attributes into the question “Is blue bird right to yellow bird?”, and the MLLMs answers differ from the ground-truth, we cannot ascertain whether the discrepancy stems from a misunderstanding of attributes (i.e., blue and yellow) or inter-object relation (i.e., right). Inspired by [44], we recognize that unusual data are crucial for evaluating and enhancing MLLM capabilities. We therefore create an adversarial relation subset comprising 710 challenging question-answer pairs. This subset is designed to rigorously assess the relation understanding capabilities of MLLMs.
In summary, MMRel is featured with four distinctive characteristics as illustrated in Fig. 3(e): (i) It comprises large-scale and diverse relation data that are collected from different sources; (ii) It comprises high-quality relation annotations and negative-choice data for action relations; (iii) It adopts an effective and efficient evaluation framework [29] that employs straightforward Yes/No instead of other complicated discriminative or generative evaluations; (iv) It includes an adversarial subset designed to evaluate the MLLMs’ ability to understand challenging relations.
4 Experiments
In this section, we systematically analyze the effectiveness of Multi-Modal Relation Understanding (MMRel). We first describe the experiment setups in Sec. 4.1. Subsequently, we employ MMRel to evaluate the relation understanding capabilities of multiple representative Multi-Modal Large Language Models (MLLMs) and hallucination mitigation techniques in Sec. 4.2. Finally, we employ MMRel to fine-tune LLaVA-1.5 [2] to examine how much the fine-tuning enhances their relation understanding capabilities, which is further discussed in Sec. 4.3.
Table 4: Evaluation results of existing MLLMs on our MMRel benchmark. Existing MLLMs can not handle the task of relation understanding well.
Domain Taxonomy Model Accuracy(%) Precision(%) Recall(%) F1-Score(%)
Real Spatial InstructBLIP [5] 50.58 50.36 81.38 62.22
LLaVA-1.5  [2] 51.48 51.01 74.88 60.68
Qwen-VL [7] 54.10 52.65 81.34 63.93
VCD [32] 50.76 50.42 92.01 65.14
DOLA [30] 51.54 51.00 78.46 61.82
OPERA [31] 51.21 50.68 90.24 64.91
Action InstructBLIP [5] 65.14 60.59 89.92 72.40
LLaVA-1.5 [2] 67.67 62.98 88.35 73.54
Qwen-VL [7] 69.87 64.04 92.94 75.83
VCD [32] 65.33 59.86 96.60 73.91
DOLA [30] 69.10 63.53 92.09 75.19
OPERA [31] 70.35 63.74 96.68 76.83
Comparative InstructBLIP [5] 56.23 54.96 69.00 61.19
LLaVA-1.5 [2] 61.70 61.56 62.31 61.93
Qwen-VL [7] 72.19 69.95 77.81 73.67
VCD [32] 62.16 60.64 69.30 64.68
DOLA [30] 60.64 60.48 61.40 60.94
OPERA [31] 74.32 75.64 71.73 73.63
SDXL Spatial InstructBLIP [5] 51.37 50.92 75.74 60.90
LLaVA-1.5 [2] 52.46 51.50 84.55 64.01
Qwen-VL [7] 55.84 54.59 69.45 61.13
VCD [32] 50.63 50.33 96.68 66.20
DOLA [30] 52.12 51.21 89.70 65.20
OPERA [31] 50.92 50.48 97.25 66.46
Comparative InstructBLIP [5] 51.23 50.95 65.85 57.45
LLaVA-1.5 [2] 59.08 58.75 60.92 59.82
Qwen-VL [7] 70.15 77.64 56.62 65.48
VCD [32] 64.31 63.17 68.62 65.78
DOLA [30] 58.46 57.49 64.92 60.98
OPERA [31] 67.85 65.34 76.00 70.27
Dall-E Spatial InstructBLIP [5] 55.29 53.06 91.74 67.23
LLaVA-1.5 [2] 53.22 51.88 88.93 65.53
Qwen-VL [7] 57.44 54.98 82.15 65.87
VCD [32] 51.65 50.86 97.85 66.93
DOLA [30] 62.98 61.39 77.56 68.53
OPERA [31] 69.23 64.68 89.92 74.24
Action InstructBLIP [5] 63.14 61.29 79.02 69.03
LLaVA-1.5 [2] 61.20 60.80 71.38 65.67
Qwen-VL [7] 62.47 62.90 67.80 65.26
VCD [32] 61.62 58.65 88.78 70.63
DOLA [30] 53.06 51.68 94.05 66.71
OPERA [31] 50.41 50.21 97.85 66.37
4.1 Experiment Setups
MLLM Baselines. We conduct experiments with six representative MLLMs [2, 5, 7, 30, 31, 32]. Three of them, namely, LLaVA-1.5 [2], InstructBLIP [5], and Qwen-VL [7], are versatile and can work on a variety of vision-language tasks. However, they are trained with image-text pairs with coarse granularity [45, 46] and thus prone to generating hallucinations [47, 48]. We therefore select another three MLLMs that were designed for hallucination mitigation. As MMRel involves inter-object relations only, we select DOLA [30], OPERA [31], and VCD [32] that focus on mitigating object hallucinations and so are more suitable for evaluating their relation understanding capabilities. Notably, the techniques selected for managing object hallucinations do not involve additional training, ensuring fair comparisons.
Evaluation. For fair benchmarking, we evaluate all six MLLMs using their 7B versions. Specifically, LLaVA-1.5 [2] and InstructBLIP [5] incorporate Vicuna 7B as their language decoder [49], while Qwen-VL adopts Qwen-7B backbone [7]. Given the impressive performance of LLaVA-1.5 [2], we also employ DOLA [30], OPERA [31], and VCD [32] in our evaluation, all based on LLaVA-1.5 [2]. As MMRel adopts a Yes/No evaluation framework, we adopt the same metrics, accuracy, precision, recall, and F1-Score, as in POPE [29]. We follow all details of evaluation as VCD [32].
4.2 Evaluations with MMRel
We first employ the 15K question-answer pairs in MMRel to evaluate how MLLMs perform when facing three categories of relations from three data domains. As Tab. 4 shows, all six MLLMs face various challenges in relation understanding. Specifically, we observe that the performance on action relations is superior to that on spatial and comparative relations in real images. This disparity arises because the coarsely-aligned training data (all are real images) predominantly features action descriptions, whereas spatial and comparative terms are often represented by more vague expressions, such as near, next to, in front of, etc. Furthermore, because the synthetic images generated by SDXL [26] and DALL-E [27] have distributions that differ significantly from the pre-training data, their performance on action relations is lower than that observed with real images. Moreover, even techniques designed to address object hallucinations, such as those reported in [30, 31, 32], continue to struggle with relation understanding. Their limited performances underscore that the primary challenges in MMRel arise from understanding relations rather than recognizing objects. The performances in Tab. 4 could serve as baselines for further training-free methods.
Table 5: Statistics of training, test and adversarial sets in MMRel.
Domain Real SDXL Dall-E Total
Taxonomy Spatial Action Comparative Total Spatial Comparative Total Spatial Action Total
[…]
4.3 Fine-Tuning with MMRel
The large scale and high diversity of MMRel also enable it to fine-tune MLLMs to enhance their relation understanding capabilities. To verify this, we split the 15K data into training and test sets, as shown in Tab. 5. Here, we only adopt LLaVA-1.5 [2] as a strong baseline due to its popularity, outstanding performance, and versatility. To maintain the generality of the MLLMs and prevent overfitting to our training data, we have integrated the MMRel training set with 260K instruction data points from LLaVA [1]. Furthermore, adhering to the instruction template used in LLaVA [1], we have expanded our Yes/No answers into various formats to enhance the diversity of the dialogues (with more details in the Appendix). We strictly follow the steps for instruction tuning from LLaVA-1.5 [2]. The statistics of our adversarial subset are shown in Tab. 5.
Table 6: Fine-tuning LLaVA-1.5 with MMRel improves the performance significantly.
Domain Taxonomy Model Accuracy(%) Precision(%) Recall(%) F1-Score(%)
Real Spatial LLaVA-1.5 [2] 51.29 50.88 74.83 60.57
Finetune with MMRel 79.66 82.45 75.34 78.74
Action LLaVA-1.5 [2] 65.08 61.68 84.97 71.47
Finetune with MMRel 81.15 81.21 82.47 81.84
Comparative LLaVA-1.5 [2] 64.04 83.33 66.18 73.77
Finetune with MMRel 76.40 79.75 92.65 85.71
SDXL Spatial LLaVA-1.5 [2] 50.63 50.39 80.63 62.02
Finetune with MMRel 95.63 96.20 95.00 95.60
Comparative LLaVA-1.5 [2] 61.24 82.89 63.00 71.59
Finetune with MMRel 77.13 77.43 99.50 87.09
Dall-E Spatial LLaVA-1.5 [2] 65.46 51.57 89.60 65.46
Finetune with MMRel 94.06 94.95 93.07 94.00
Action LLaVA-1.5 [2] 57.69 57.94 66.83 62.07
Finetune with MMRel 71.79 67.04 89.60 76.69
Adversarial Mix LLaVA-1.5 [2] 50.35 47.10 70.50 56.47
Finetune with MMRel 67.52 60.79 81.37 69.59
Following instruction tuning with the MMRel training set, we observe significant and consistent improvements in performance across all three domains and relation categories, as detailed in Tab. 6. Additionally, such fine-tuning enhances the performance on the adversarial subset as well. To validate our hypothesis that integrating our MMRel with LLaVA instruction data [1] is beneficial to other perceptual tasks, we conduct comparative experiments using the MME benchmark [24] against the baseline LLaVA-1.5 model [2]. Experiments in Tab. 7 show that instruction tuning with MMRel achieves clear improvements for most of the MME tasks, verifying the scalability and applicability of such fine-tuning in various multi-modal tasks. The evaluation metrics of MME [24] are described in the Appendix.
Table 7: Fine-tuning LLaVA-1.5 with MMRel improves performances clearly on MME.
Method Position Existence Count Color Posters Celebrity Scene Landmark Artwork OCR Total Score
LLaVA-1.5 [2] 114.0 175.7 124.7 151.0 127.8 113.6 148.3 130.0 102.2 92.0 1279.3
FT 126.7 180.0 128.3 153.3 129.6 122.1 149.3 131.9 117.0 117.5 1355.7
5 Conclusion
We notice that Multi-Modal Large Language Models (MLLMs) encounter difficulties in understanding inter-object relations, a research area that remains under-explored. To address this gap, in this paper, we first provide a clear taxonomy and define relation hallucinations for inter-object relations, and further introduce the Multi-Modal Relation Understanding (MMRel) dataset. MMRel is specifically designed to probe inter-object relations in MLLMs and includes a challenging adversarial subset featuring highly unusual relations. Thanks to its large-scale, high-quality and diverse data, MMRel serves as a versatile tool for evaluating and fine-tuning MLLMs, enhancing their capabilities in relation understanding and other vision-language tasks. Extensive experiments across various popular MLLMs validate MMRel’s effectiveness.
Acknowledgments and Disclosure of Funding
This study is supported by the Interdisciplinary Graduate Programme, Nanyang Technological University. We thank students from Nanyang Technological University: Zhixiao Jin, Siwei Huang, Guanqiao Fu, Yu Wei, and Taorui Wang for establishing the MMRel dataset.
Appendix
Appendix A Limitations and Future Work
Despite conducting extensive experiments to verify the effectiveness of MMRel, our fine-tuning efforts are currently limited to using LLaVA-1.5 [2] as the base model. Besides, we only adopt Yes/No evaluation just like POPE [29]. Future work should consider incorporating generative evaluation metrics, such as those used in CHAIR [47], to assess object hallucinations more comprehensively. Furthermore, the dataset for comparative relations remains relatively limited, thus, there is a need to collect and annotate more data within this subset. Lastly, the taxonomy of inter-object relations could be expanded to include additional categories.
Appendix B MMRel Dataset
Sample Images. We specifically aim to generate images featuring unusual inter-object relations in multiple styles using DALL-E-3 [27]. Additional examples of synthetic images from Dall-E-3 [27] are shown in Fig. A1.
Figure A1: Illustration of sample images generated by Dall-E. These images contain unusual relations with different styles.
SemiDC via GPT-4V [28]. We employ an in-context learning paradigm to label images in the Visual Genome (VG) dataset [12] using GPT-4V [28]. The specific prompts used are illustrated in Fig. A2
Figure A2: In-context learning prompts for labeling images from VG [12].
SemiDC via Dall-E-3 [27]. We use the following prompts as inputs for DALL-E 3 [27] to generate images: a style photo depicting a strange or unusual object-object relation: inter-object relations. Here, style specifies the image style, selected from photo-realistic, watercolor, abstract, and oil painting, and inter-object relations describes the specific relation we aim to depict. For example, the prompt for generating the first image in Fig. A1 is: an abstract photo depicting a strange or unusual object-object relation: cat on toilet.
Appendix C Experiments
Data for Fine-tuning LLaVA-1.5 [2]. We integrate the training set of MMRel with 260K original instances from the total 665K dataset utilized for LLaVA [2] instruction tuning to maintain the versatility of MLLMs and prevent overfitting to our specialized training data. Specifically, this 260K subset includes question-answer instructions comprising 73K data points from COCO [50], 53K from VG [12], 72K from GQA [51], 22K from TextVQA [52], and 41K text-only data.
Question-answer Instructions. The original annotations are in the form of triplets, such as “cat on car”. We expand these into both discriminative and generative question-answer pairs.
The discriminative instructions are as follows:
Positive:
Q: Is there a cat on a car in the image? Please answer with one word.
A: Yes.
Negative:
Q: Is there a car on a cat in the image? Please answer with one word.
A: No.
The generative instructions are as follows:
Q: What is the spatial relation between a cat and a car in the image?
A: A cat is on a car.
Evaluation on MME [24]. MME [24] employs a Yes/No evaluation framework. Adhering to the original settings, we calculate the final score using the sum of accuracy and accuracy+, where accuracy is determined for each individual question and accuracy+ is assessed for each image, requiring correct answers to both associated questions. Accuracy+ serves as a stricter metric, offering a more comprehensive reflection of the MLLMs’ capabilities.