### v2

# Jiahao Nie1,2    Gongjie Zhang311footnotemark: 1    Wenbin An4  Yap-Peng Tan2  Alex C. Kot2  Shijian Lu2

Abstract
Though Multi-modal Large Language Models (MLLMs) have recently achieved significant progress, they often face various problems while handling inter-object relations, i.e., the interaction or association among distinct objects. This constraint largely stems from insufficient training and evaluation data for relation understanding, which has greatly impeded MLLMs in various vision-language generation and reasoning tasks. We attempt to address this challenge by introducing Multi-Modal Relation Understanding (MMRel), a benchmark that features large-scale, high-quality, and diverse data on inter-object relations. MMRel features three distinctive attributes: (i) It contains over 22K question-answer pairs, spanning three distinct domains and covering three relation categories, ensuring both scale and diversity; (ii) it provides manually verified, high-quality labels to ensure exceptional annotation accuracy; (iii) it includes adversarial cases with highly unusual relations, offering a challenging setting for evaluating relation hallucination. These features make MMRel ideal for evaluating MLLMs on relation understanding, as well as for fine-tuning MLLMs to enhance relation comprehension capability. Extensive experiments verify the effectiveness of MMRel in evaluating and enhancing MLLMs’ relation understanding capabilities. The benchmark has been released publicly.
1 Introduction
Multi-Modal Large Language Models (MLLMs) [40, 39, 63, 30, 18, 22, 4, 64] have demonstrated remarkable capabilities in various vision-language tasks [43, 44, 66, 65, 31, 42, 13, 2, 61, 14, 11, 59, 47]. One essential component while handling these tasks is to understand the complicated relation among various objects in images [70] (e.g., the interaction or association between distinct objects [27]), which is crucial for advanced reasoning accurate VQA [28, 55, 3] and detailed image captioning [33, 58]. However, recent studies [70, 19, 41, 6] reveal that most existing MLLMs tend to struggle in interpreting inter-object relations, as illustrated in Fig. 1. This challenge arises primarily from lacking a large-scale, diverse, and high-quality benchmark tailored for MLLMs in the current era.
Figure 1: MLLMs often fail to understand inter-object relations.
Table 1: A comparison of MMRel with existing relation understanding benchmarks. MMRel is featured with large-scale, diverse, and high-quality data on inter-object relations.
Benchmark #Q&A Taxonomy Domain MLLM+++ Adv. Dis. Gen.
Spatial Action Comparative Real SDXL Dall-E Human Subset QA QA
MMHAL [52] 12 $\checkmark$ $\times$× $\times$× $\checkmark$ $\times$× $\times$× $\checkmark$ $\times$× $\times$× $\checkmark$
RAH [15] 500 $\checkmark$ $\checkmark$ $\times$× $\checkmark$ $\times$× $\times$× $\times$× $\times$× $\checkmark$ $\times$×
MERLIM [54] - $\checkmark$ $\times$× $\times$× $\checkmark$ $\times$× $\times$× $\times$× $\times$× $\checkmark$ $\times$×
FAITHSCORE [25] - - - - $\checkmark$ $\times$× $\times$× $\checkmark$ $\times$× $\times$× $\checkmark$
M-HalDetect [23] - - - - $\checkmark$ $\times$× $\times$× $\checkmark$ $\times$× $\times$× $\checkmark$
AMBER [56] 1.7K $\checkmark$ $\times$× $\times$× $\checkmark$ $\times$× $\times$× $\checkmark$ $\times$× $\checkmark$ $\times$×
SPEC [45] 3.5K $\checkmark$ $\times$× $\checkmark$ $\times$× $\checkmark$ $\times$× $\times$× $\times$× $\checkmark$ $\times$×
Hallusion [37] 150 $\times$× $\times$× $\checkmark$ $\checkmark$ $\times$× $\times$× $\times$× $\times$× $\checkmark$ $\times$×
MME [45] 60 $\checkmark$ $\times$× $\times$× $\checkmark$ $\times$× $\times$× $\times$× $\times$× $\checkmark$ $\times$×
R-Bench [60] 11.7K $\checkmark$ $\checkmark$ $\times$× $\checkmark$ $\times$× $\times$× $\checkmark$ $\times$× $\checkmark$ $\times$×
MMRel 22.5K $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Though several benchmarks on inter-object relations have been created, they were not intended for assessing MLLMs’ relation understanding capabilities [56, 52, 45, 23, 15, 54, 25, 21, 37, 60]. Specifically, most existing benchmarks suffer from obvious limitations in data scales, relation categories, and data diversity as shown in Table 1. We address this issue by creating a comprehensive benchmark on inter-object relations, aiming to gauge and enhance MLLMs’ relation understanding capability in various multimodal tasks. Specifically, we first taxonomize inter-object relations by three distinct categories, namely, spatial relations (e.g., a dog is left to a cat), action relations (e.g., a boy eats a burger), and comparative relations (e.g., an apple is smaller than a watermelon). Additionally, we introduce relation hallucination, a new concept referring to the scenario when MLLM-generated relations align with the common sense but contradict the actual contents in images (e.g., a MLLM responds by “a man drives a car” but the actual image content is “a man pushes a car”.
The created Multi-Modal Relation Understanding benchmark (MMRel) has several desirable features. First, it is large-scale comprising multi-modal data of three types of inter-object relations (i.e., spatial, action, and comparative) as well as a challenging adversarial subset that consists of abnormal relations deviating from the common sense. All these lead to around 22,500 Q&A pairs, each of which comes with clearly defined annotations and relation categories. Second, it has great diversity where the image data were collected from three distinctive sources including real images, SDXL [46] generated images with clean background, and Dall-E [7] generated multi-style images as illustrated in Fig. 2. Third, it comes with high-quality data as collected via a semi-automatic pipeline which first generates images and annotations with MLLMs [7, 1] and then reviews and corrects them by human annotators. Last but not least, it consists of both the widely adopted discriminative questions with Yes/No answers [56, 15, 54, 21, 37, 35, 60, 71] and generative open-ended questions via LLM-assisted evaluations [40, 52].
Figure 2: Sample images from MMRel. MMRel includes three categories of inter-object relations, spatial, action, and comparative. The images are sourced from three domains: (a) real images, (b) synthetic images generated by SDXL, and (c) synthetic images generated by Dall-E. Additional Dall-E samples are included in the supplementary material.
MMRel can serve different purposes thanks to its large-scale, diverse, and high-quality multimodal data. The primary usage is to gauge existing MLLMs in relation understanding. Extensive evaluations on MMRel reveal that most general MLLMs [39, 18, 4, 62, 9], grounding-specific MLLMs [12], and even hallucination mitigation methods [17, 24, 29] exhibit sub-optimal performance in understanding inter-object relations. Another typical usage is MLLM fine-tuning for enhanced relation understanding capability [40, 39]. Our experiments show that MMRel as training data can clearly improve MLLMs’ relation understanding and related perception capabilities.
The contributions of this work can be summarized in three aspects. First, we systematically define inter-object relations and relation hallucinations under the context of MLLMs, laying a solid foundation for future study of relation hallucinations and their mitigation in MLLMs. Second, we contribute MMRel, a large-scale benchmark with over 22K high-quality vision-language samples sourced from diverse domains, each with clearly defined inter-object relations. To the best of our knowledge, this is the first large-scale benchmark dedicated to inter-object relations with superior diversity and quality. Third, extensive experiments demonstrate MMRel’s effectiveness for both evaluating current MLLMs and fine-tuning them to enhance relation understanding capabilities.
2 Existing Benchmarks on Relation Understanding
The study of inter-object relations boasts a rich history, with Scene Graph Generation (SGG) [26, 68, 53] serving as a foundational area of focus. SGG was collected for identifying objects and their interrelations within images. Leveraging specialized object detection methods [49, 36] that excel in detecting both tiny and inconspicuous objects, SGG benchmarks such as Visual Genome (VG) [27] are distinguished by their extremely detailed and comprehensive annotations. Following their recent successes [48], Multi-Modal Large Language Models (MLLMs) have received significant attention. This has inspired several new benchmarks [19, 67] which were collected for assessing the relation understanding capabilities of MLLMs.
Specifically, VL-Checklist [70] and ARO [67] are two representative benchmarks that evaluate MLLMs’ capabilities of understanding fine-grained information, including inter-object relations. The two benchmarks inherit images and labels from VG [27] and adopt a binary Positive/Negative evaluation framework, as illustrated in Fig. 3(a)-(b). However, both benchmarks have significant limitations, as they generate negative choices randomly without human verification. Specifically, they generate negative choices by: (i) randomly altering the relations (e.g., man riding shirt); (ii) simply reversing the order of objects (e.g., grass is eating the horse), making the generated negative options semantically implausible or even impossible. With such designs, MLLMs can easily deduce the correct answers even without seeing images, as the powerful LLM enables MLLMs to effectively dismiss those negative choices that starkly deviate from the common sense.
Moreover, although recent MLLMs [40, 39, 63, 30, 18, 22, 4] are versatile, their proficiency in object detection is limited due to pre-training on coarse granularity image-text pairs [51, 10]. Therefore, the overly dense objects, occasionally incorrect annotations, and unclear taxonomies from existing benchmarks, such as VG [27], are unsuitable for assessing MLLM. Consequently, re-labeling some existing benchmarks is essential for developing new benchmarks that are more suitable for evaluating relation understanding in the MLLM era. This has triggered recent initiatives such as M-HalDetect [23] and AMBER [56] that concentrate on generating new labels. Specifically, M-HalDetect [23] incorporates manually annotated relations within lengthy descriptive sentences, with evaluations performed by human assessors (refer to Fig. 3(c)). However, this metric does not straightforwardly reflect the capabilities in relation understanding and introduces potential risks of subjective comparisons. On the other hand, AMBER [56] simplifies the classification of relationships to contact or not and employs a Yes/No evaluation framework (refer to Fig. 3(d)). While this approach is straightforward, it leads to an imprecise evaluation due to the restricted definition of relations.
The limitations of existing benchmarks can be summarized as: (i) They are small-scale and lack diversity; (ii) Their annotations are of low quality, with negative choices for action relations often being meaningless; (iii) They lack clear definitions and taxonomies of inter-object relations.
Figure 3: Comparison of MMRel with existing benchmarks. (a)-(b) negative choices for action relations are implausible and can be easily ruled out; (c) evaluation metrics are overly complex and subjective; (d) considering only one relation “contact”, which is insufficient and ambiguous. MMRel mitigates these limitations by: (i) providing a comprehensive taxonomy of inter-object relations, (ii) carefully designing plausible negative action relations, and (iii) adopting discriminative Yes/No and generative open-ended evaluations.
3 Multi-Modal Relation Understanding Benchmark (MMRel)
As analyzed in Sec. 1 and Sec. 2, existing benchmarks for relation understanding are generally constrained by their small-scale and homogeneous nature. This motivates us to develop a new benchmark, Multi-Modal Relation Understanding (MMRel), characterized by large-scale, diverse, and high-quality data on inter-object relations. In this section, we first describe the image sources and relation taxonomies in Sec. 3.1. Subsequently, we introduce a Semi-automatic Data Collection pipeline in Sec. 3.2, designed to amass large-scale, diverse, and high-quality inter-object relation data. Finally, we present the statistics and additional details of the MMRel in Sec. 3.3.
3.1 Data Source and Relation Taxonomy
Following previous works [70, 67], we source our real image subset from the Visual Genome (VG) [27] and select synthetic images from the SPEC [45], which is based on SDXL [46]. Specifically, we use the relative spatial subset for spatial relations and the relative size subset for comparative relations. To further diversify the MMRel, we also generate synthetic images via Dall-E [7]. Given the good clarity of annotations in SPEC [45], we adapt their labels to enhance our benchmark’s utility for relation understanding. Details regarding annotation and image generation across other domains are outlined in Sec. 3.2.
Moreover, we employ distinct annotation strategies for each of three relation taxonomies. For spatial relations, we limit the categories to left, right, on/over, and under, as other descriptors such as around or next to do not accurately capture relative inter-object positions. For action relations, our annotations are based on the capabilities of Multi-Modal Large Language Models (MLLMs), as detailed in Sec. 3.2. Comparative relations are annotated manually. Our provided taxonomies of relations are more suitable in the MLLM era, which is further discussed in the supplementary material. We also analyze the complexity of MMRel in the supplementary material.
3.2 Semi-Automatic Pipeline
In this section, we introduce our Semi-automatic Data Collection pipeline (SemiDC), as illustrated in Fig. 4. This pipeline is capable of annotating large-scale existing images and generating a substantial amount of high-quality synthetic images. As discussed in Sec. 2, re-labeling existing images is essential since their original labels are incompatible with MLLMs. To this end, we design SemiDC to generate high-quality relation annotations via GPT-4V [1] (refer to Fig. 4(a)) for large-scale VG benchmark [27]. This process is divided into three stages: (i) Pre-processing: We selectively exclude images featuring complex scenes that pose challenges for GPT-4V [1] in generating accurate annotations. Specifically, we utilize the object labels and corresponding bounding boxes from the VG benchmark [27] to identify and exclude images that contain more than 10 tiny objects, each occupying less than 1/20 of the total image area; (ii) Re-labeling via GPT-4V [1]: We employ the in-context learning paradigm [30, 69] to use GPT-4V [1] to generate relation annotations. For example, we provide an image and its accurate relation annotations as a reference, together with a query image accompanied by the textual prompt “Please generate the action relationships between salient objects in this image”. GPT-4V [1] will then produce annotations for the query image based on the text prompt. More details can be found in the supplementary material; (iii) Human verification: We manually assess and correct the annotations that are generated by GPT-4V [1], to ensure the quality of the collected inter-object relation data.
Figure 4: Pipeline of Semi-automatic Data Collection (SemiDC). (a) Re-labeling real images from the Visual Genome using GPT-4V, with human verification for correctness. (b) Generating synthetic images based on textual prompts via Dall-E, followed by human verification to ensure accuracy.
Table 2: Statistics of MMRel, left: statistics of each subset, right: properties of the complete MMRel.
Domain Real SDXL Dall-E Total
Taxonomy Spat. Act. Comp. Total Spat. Comp. Total Spat. Act. Total
#Y/N QA 4.86K 4.62K 0.66K 10.14K 1.75K 0.65K 2.40K 1.20K 1.18K 2.39K 14.93K
#Gen. QA 2.43K 2.35K 0.32K 5.10K 0.87K 0.33K 1.20K 0.61K 0.62K 1.22K 7.51K
Complete MMRel
Property Obj. Spat. Act. Comp.
Number 804 5 380 9
As MLLMs are known to have a bias towards positive responses [35, 38], we design MMRel to maintain a balance between “Yes” and “No” answers to mitigate potential biases in evaluation metrics [35] for binary evaluation. Specifically, negative for spatial and comparative relations can be simply derived by reversing the order of two objects. For instance, “dog is left to cat” is the inverse of “cat is left to dog”. Nevertheless, as discussed in Sec. 2, randomly generated negative choices for action relations are implausible, which can be easily ruled out even without referring to the images. To address this issue, we first adopt GPT-3.5 [8] to generate plausible negative relations that are well aligned with common sense for given objects, and then verify the generated negative relations by manual checking. For instance, a suitable negative relation for “man eats cake” might be generated as “man bakes cake”. This design significantly improves over prior works [70, 67] in terms of both training and evaluation. In addition, we provide a potential to leverage LLM to evaluate MLLMs’ capabilities of relation understanding in open-ended questions, such as “What is the action relation between a dog and a frisbee?”.
Furthermore, we generate more diverse inter-object relation data via generative models. The SemiDC pipeline for generating images (refer to Fig. 2(b)) is also divided into three stages: (i) Preparing textual prompts: We prepare textual prompts describing inter-object relations; (ii) Image generation via Dall-E [7]: We utilize the prepared text prompts as inputs for Dall-E-3 [7] to generate corresponding images. For instance, if we want Dall-E [7] to generate an image depicting a crocodile on the desert, then we should employ the prompt “a photo-realistic picture depicting an object-object relation: crocodile on desert” as input. These textual prompts define our targeted relations and the visual style of the generated images. We consider four styles, photo-realistic, watercolor, abstract, and oil painting (refer to Fig. 2(c)), to enrich our benchmark; (iii) Human verification of generated images: If the generated images do not match the input relations, we will modify the relations corresponding to generated images. The same as annotating negative choices for VG [27], we utilize GPT-3.5 [8] to generate negative action relations. More details of SemiDC for both relabeling and generation of inter-object relation data are described in the supplementary material.
3.3 Statistics of MMRel
Tab. 2 shows the statistics of MMRel. Specifically, MMRel comprises around 22,500 question-answer pairs (15K Yes/No, and 7.5K Open-ended) across 7 subsets, spanning 3 domains and 3 categories of relations. Concretely, approximately 68% of the data consists of real images, while the remaining 32% comprises synthetic images (i.e., SDXL [46] and Dall-E [7]). In terms of relation categories, MMRel includes a larger scale of spatial and action relations compared to comparative relations. Thanks to the open-vocabulary capability of GPT-4V [1], MMRel guarantees a diverse range of objects and action relations, as shown in Tab. 2. Regarding comparative relations, we manually re-label VG images, resulting in a subset with accurate annotations. It is important to note that our annotations strictly describe the objects and their relations, without including any attributes, so that the evaluation can accurately reflect the relation understanding capabilities. For example, if we introduce attributes into the question “Is blue bird right to yellow bird?”, and the MLLMs answers differ from the ground-truth, we cannot ascertain whether the discrepancy stems from a misunderstanding of attributes (i.e., blue and yellow) or inter-object relation (i.e., right). Inspired by [5], we recognize that unusual data are crucial for evaluating and enhancing MLLM capabilities. We therefore create an adversarial relation subset comprising 710 challenging question-answer pairs. These pairs are meticulously manually selected from the complete MMRel to include images from all domains and encompass all relation categories. This subset is designed to rigorously assess the relation understanding capabilities of MLLMs, the statistics of which are described in Sec. 4.
Table 3: Evaluation results of various MLLMs on the MMRel. Results indicate that these MLLMs (including proprietary models and grounding-specific models) still struggle with inter-object relation understanding.
Domain Taxonomy Model Accuracy(%) Precision(%) Recall(%) F1-Score(%)
Real Spatial InstructBLIP [18] 50.58 50.36 81.38 62.22
LLaVA-1.5  [39] 50.45 50.30 75.04 60.23
Qwen-VL [4] 54.10 52.65 81.34 63.93
MiniCPM-V [62] 52.88 51.85 80.64 63.12
Shikra [12] 50.89 50.54 82.00 62.54
VCD [29] 50.76 50.42 92.01 65.14
DOLA [17] 51.54 51.00 78.46 61.82
OPERA [24] 51.21 50.68 90.24 64.91
GPT-4o [9] 59.08 55.10 98.11 70.57
Action InstructBLIP [18] 65.14 60.59 89.92 72.40
LLaVA-1.5 [39] 67.67 62.98 88.35 73.54
Qwen-VL [4] 69.87 64.04 92.94 75.83
MiniCPM-V [62] 76.54 70.06 94.05 80.30
Shikra [12] 70.84 65.02 92.34 76.31
VCD [29] 65.33 59.86 96.60 73.91
DOLA [17] 69.10 63.53 92.09 75.19
OPERA [24] 70.35 63.74 96.68 76.83
GPT-4o [9] 79.69 71.91 98.55 83.15
Comparative InstructBLIP [18] 56.23 54.96 69.00 61.19
LLaVA-1.5 [39] 61.70 61.56 62.31 61.93
Qwen-VL [4] 72.19 69.95 77.81 73.67
MiniCPM-V [62] 62.46 59.53 77.81 67.46
Shikra [12] 62.16 58.97 79.94 67.87
VCD [29] 62.16 60.64 69.30 64.68
DOLA [17] 60.64 60.48 61.40 60.94
OPERA [24] 74.32 75.64 71.73 73.63
GPT-4o [9] 94.83 93.26 96.66 94.93
SDXL Spatial InstructBLIP [18] 51.37 50.92 75.74 60.90
LLaVA-1.5 [39] 51.37 50.87 80.66 62.39
Qwen-VL [4] 55.84 54.59 69.45 61.13
MiniCPM-V [62] 52.40 51.55 80.10 62.72
Shikra [12] 51.43 50.83 87.41 64.28
VCD [29] 50.63 50.33 96.68 66.20
DOLA [17] 52.12 51.21 89.70 65.20
OPERA [24] 50.92 50.48 97.25 66.46
GPT-4o [9] 79.69 71.91 98.55 83.15
Comparative InstructBLIP [18] 51.23 50.95 65.85 57.45
LLaVA-1.5 [39] 59.08 58.75 60.92 59.82
Qwen-VL [4] 70.15 77.64 56.62 65.48
MiniCPM-V [62] 58.46 55.77 81.85 66.33
Shikra [12] 64.31 61.10 78.77 68.82
VCD [29] 64.31 63.17 68.62 65.78
DOLA [17] 58.46 57.49 64.92 60.98
OPERA [24] 67.85 65.34 76.00 70.27
GPT-4o [9] 71.00 74.58 68.62 71.47
Dall-E Spatial InstructBLIP [18] 55.29 53.06 91.74 67.23
LLaVA-1.5 [39] 52.98 51.71 90.08 65.70
Qwen-VL [4] 57.44 54.98 82.15 65.87
MiniCPM-V [62] 60.41 56.02 96.86 70.99
Shikra [12] 49.42 49.68 91.07 64.29
VCD [29] 51.65 50.86 97.85 66.93
DOLA [17] 62.98 61.39 77.56 68.53
OPERA [24] 69.23 64.68 89.92 74.24
GPT-4o [9] 75.21 67.15 98.68 79.92
Action InstructBLIP [18] 63.14 61.29 79.02 69.03
LLaVA-1.5 [39] 60.27 60.28 69.11 64.39
Qwen-VL [4] 62.47 62.90 67.80 65.26
MiniCPM-V [62] 70.92 75.71 64.88 69.88
Shikra [12] 66.19 63.69 81.30 71.43
VCD [29] 61.62 58.65 88.78 70.63
DOLA [17] 53.06 51.68 94.05 66.71
OPERA [24] 50.41 50.21 97.85 66.37
GPT-4o [9] 81.07 78.13 88.29 82.90
Table 4: Statistics of training, test and adversarial (a challenging subset for testing) splits when MMRel serves as fine-tuning data, upper: number of images in each split, lower: number of distinct object and relation in each split.
Domain Real SDXL Dall-E Total
Taxonomy Spatial Action Comparative Total Spatial Comparative Total Spatial Action Total
[…]
Domain Real SDXL Dall-E
Taxonomy Spatial Action Comparative Spatial Comparative Spatial Action
Property Object Relation Object Relation Object Relation Object Relation Object Relation Object Relation Object Relation
training split 475 5 111 304 149 9 75 4 40 2 164 5 151 141
test split 226 5 111 304 39 9 52 4 28 2 133 5 136 141
In summary, MMRel is featured with four distinctive characteristics as illustrated in Fig. 3(e): (i) It comprises large-scale and diverse relation data that are collected from different sources; (ii) It comprises high-quality relation annotations and negative-choice data for action relations; (iii) It adopts an effective and efficient evaluation framework [35] that employs straightforward discriminative Yes/No, and it provides a potential to evaluate generative open-ended questions; (iv) It includes an adversarial subset designed to assess the MLLMs’ ability to understand challenging relations.
4 Experiments
In this section, we systematically analyze the effectiveness of Multi-Modal Relation Understanding benchmark (MMRel). We first describe the experiment setups in Sec. 4.1. Subsequently, we employ MMRel to evaluate the relation understanding capabilities of multiple representative Multi-Modal Large Language Models (MLLMs) and hallucination mitigation techniques in Sec. 4.2. Finally, we employ MMRel to fine-tune LLaVA-1.5 [39] to examine how much the fine-tuning enhances their relation understanding capabilities, which is further discussed in Sec. 4.4.
4.1 Experiment Setups
MLLM Baselines. We evaluate MMRel over nine representative MLLMs [39, 18, 4, 12, 62, 17, 24, 29, 9]. Specifically, four MLLMs including LLaVA-1.5 [39], InstructBLIP [18], Qwen-VL [4], and MiniCPM-V [62] are open-sourced and capable of a variety of vision-language tasks. Nevertheless, they are trained with image-text pairs of coarse granularity [34, 32] and thus prone to generating hallucinations [50, 57]. Another three MLLMs were trained for hallucination mitigation. Since MMRel focuses on inter-object relations, we select DOLA [17], OPERA [24], and VCD [29] that aim for mitigating object hallucinations and are therefore suitable for evaluations on inter-object relations. In addition, we select another two MLLMs, namely, Shikra [12] that is trained for fine-grained grounding tasks and the latest proprietary GPT-4o [9] that demonstrates amazing capabilities on a bunch of vision-language tasks.
Evaluation. For benchmarking fairness, we evaluate all open-sourced MLLMs using their 7B (or similar scale) versions. Specifically, LLaVA-1.5 [39], InstructBLIP [18], and Shikra [12] incorporate Vicuna 7B as their language decoder [16]. MiniCPM-V [62] selects LLaMA-3 [20] and Qwen-VL adopts Qwen-7B backbone [4]. The three hallucination mitigation models including DOLA [17], OPERA [24], and VCD [29] are all based on LLaVA-1.5 [39]. For discriminative Yes/No evaluation framework, we adopt the same metrics, accuracy, precision, recall, and F1-Score, as in POPE [35]. For generative open-ended questions, our evaluations are LLM-assisted. We follow all details of evaluation as VCD [29].
4.2 MLLM Evaluations with MMRel
We employ all of the 15K Yes/No question-answer pairs in MMRel to evaluate how MLLMs perform while handling multimodal data with rich inter-object relations. As Tab. 3 shows, all nine MLLMs face various problems while handling relation understanding. Several points can be further drawn from the experimental results.
Regarding the types of inter-object relations, we can observe that MLLMs perform clearly better on action and comparative relations for real images. The better performance can be largely attributed to the coarsely-aligned MLLM training data (all are real images) that predominantly capture action and comparative descriptions instead of spatial terms which often appear with vague expressions such as near, next to, in front of, etc. From the perspective of synthetic vs. real data, we can see that MLLMs generally perform better on real images as compared with synthetic images by SDXL [46] (on comparative relations) and DALL-E [7] (on action relations). The better performance is largely attributed to the distribution gap between the two sets of synthetic images and the real images that were used in MLLM training.
Table 5: LLM-assisted open-ended evaluation of some MLLMs on our MMRel benchmark. The full mark is 10.
Domain Taxonomy Model Avg. Score
Real Spatial InstructBLIP [18] 1.27
LLaVA-1.5 [39] 1.31
Qwen-VL [4] 1.22
Action InstructBLIP [18] 5.28
LLaVA-1.5 [39] 5.26
Qwen-VL [4] 5.28
Comparative InstructBLIP [18] 5.05
LLaVA-1.5 [39] 5.12
Qwen-VL [4] 5.27
SDXL Spatial InstructBLIP [18] 1.23
LLaVA-1.5 [39] 1.20
Qwen-VL [4] 1.14
Comparative InstructBLIP [18] 5.50
LLaVA-1.5 [39] 5.05
Qwen-VL [4] 5.30
Dall-E Spatial InstructBLIP [18] 2.03
LLaVA-1.5 [39] 1.97
Qwen-VL [4] 1.96
Action InstructBLIP [18] 4.82
LLaVA-1.5 [39] 4.79
Qwen-VL [4] 4.80
With respect to open-sourced vs. proprietary MLLMs, we notice that GPT-4o [9] performs much better than all open-sourced MLLMs [39, 18, 4, 12, 62]. This demonstrates the critical role of large-scale data and large models in relation understanding, underscoring the great potential for further advancements in open-sourced MLLMs.
Regarding the recent works on hallucination mitigation, we observe that all three tested methods [17, 24, 29] remain problematic in relation understanding, indicating that the primary challenges in MMRel arise from relation understanding rather than object recognition.
Note the experiments in Tab. 3 can serve as baselines for future investigation on training-free methods.
4.3 LLM-assisted Open-ended Evaluation
Motivation. Even though discriminative Yes/No evaluation is widely adopted for its straightforward merit, evaluating generative open-ended answers is gaining traction because it more closely resembles natural conversations or dialogues. Thereafter, we additionally introduce an LLM-assisted open-ended evaluation for relation understanding, which constructs MMRel to be more comprehensive. More details about question-answer pairs are discussed in the supplementary materials.
Table 6: Fine-tuning LLaVA-1.5 with MMRel significantly improves relation understanding performance.
Domain Tax. Model Acc.(%) Prec.(%) Rec.(%) F1(%) #FP
Real Spat. LLaVA [39] 51.29 50.88 74.83 60.57 419
Finetune 79.66 82.45 75.34 78.74 93
Act. LLaVA [39] 65.08 61.68 84.97 71.47 274
Finetune 81.15 81.21 82.47 81.84 99
Comp. LLaVA [39] 64.04 83.33 66.18 73.77 9
Finetune 76.40 79.75 92.65 85.71 16
SDXL Spat. LLaVA [39] 50.63 50.39 80.63 62.02 127
Finetune 95.63 96.20 95.00 95.60 6
Comp. LLaVA [39] 61.24 82.89 63.00 71.59 26
Finetune 77.13 77.43 99.50 87.09 58
Dall-E Spat. LLaVA [39] 52.72 51.57 89.60 65.46 170
Finetune 94.06 94.95 93.07 94.00 10
Act. LLaVA [39] 57.69 57.94 66.83 62.07 98
Finetune 71.79 67.04 89.60 76.69 89
Adv. Mix LLaVA [39] 50.35 47.10 70.50 56.47 255
Finetune 67.52 60.79 81.37 69.59 169
LLM-assisted Evaluation. We adopt GPT-4o [9] to score the responses from MLLMs (range 0 to 10), which is similar to LLaVA-Bench [40]. The specific prompts are as in the supplementary materials.
Experiments and Analysis. As discussed in Sec. 4, methods for grounding tasks [12] and hallucination mitigation [29, 24, 17] still struggle with relation understanding, thus we evaluate the performances of three representative general-purpose MLLMs via open-ended questions. From the results in Tab. 5, we obtain the following conclusions: First, spatial relations are more challenging than action and comparative relation, which may caused by that training data describe coarse inter-object spatial relationships (i.e., near, behind, next to) instead of specific ones in MMRel (i.e., left, right, on, under). Second, action relations in the Dall-E domain pose more difficulties than action in the real domain because the data distribution of synthetic images differs from that of training data. Third, MLLMs perform not well on inter-object relation understanding. These observations are the same as those in Yes/No evaluations in Sec. 4, which demonstrate the effectiveness and fairness of the LLM-assisted open-ended evaluation. Nevertheless, the performance differences between MLLMs are relatively small, which cannot obviously reflect the comparative capabilities between them. We assume this limitation may come from (i) The LLM used for evaluation, GPT-4o [9], is not oracle for relation understanding, which is discussed in Sec. 4; (ii) LLM-assisted evaluations and open-ended questions are subjective and less straightforward compared to the clarity and simplicity of Yes/No evaluations. It is noteworthy that our proposed LLM-assisted evaluation provides a potential and reference for open-ended questions, which is better than only including Yes/No evaluations. We believe that open-ended questions are trends but are confined by the capabilities of current LLMs.
4.4 Fine-tuning with MMRel
Data Splits. The large-scale and diverse relation data in MMRel enable it to fine-tune MLLMs for enhanced relation understanding. We verify this by splitting the 15K Yes/No question-answer pairs into a training set and a test set that consists of 11K and 4K question-answer pairs, respectively. Table 4 shows the statistics of the split data. In addition, we also collect a subset of question-answer pairs of adversarial relations as discussed in Sec. 1 and 3. The subset was manually picked from data of different domains with different types of relations, aiming to gauge MLLMs’ relation understanding under challenging scenarios. The lower part of Tab. 4 shows this adversarial subset.
Setup. We adopt LLaVA-1.5 [39] as the baseline due to its popularity, outstanding performance, and versatility. While conducting the fine-tuning, we combine the MMRel training data with 260K instruction data from LLaVA-1.5 [40] to maintain the generality of the MLLMs and prevent overfitting to our training data. Following the instruction templates used in LLaVA [40], we expanded our Yes/No answers into various formats to enhance the diversity of the dialogues (more details are in the supplementary material). We strictly followed the instruction tuning steps from LLaVA-1.5 [39] as well.
Figure 5: Fine-tuning LLaVA-1.5 with MMRel improves performances clearly on MME, validating MMRel benefits other tasks.
Results. As Tab. 6 shows, fine-tuning with MMRel improves the capabilities of relation understanding significantly and consistently across all data domains and relation categories. In addition, fine-tuning improves the relation understanding of the adversarial subset as well. Specifically, the narrow definition of hallucinations refers to the scenario when MLLMs answer “yes” while the ground-truth is “no” or vice versa. The reductions in “False Positive (FP)” across most subsets in Tab. 6 indicate that fine-tuning with MMRel suppresses hallucinations effectively. Moreover, the fine-tuning with MMRel also mitigates the MLLMs’ bias toward answering “yes” rather than “no” [35, 38]. For instance, the “yes” ratio (for Real images) for spatial and action relations decreases to 45.68% and 52.28%, respectively, from the original 73.53% and 70.93% without the fine-tuning.
Discussion. To verify that fine-tuning MLLMs like LLaVA [40] with MMRel is beneficial to other perception tasks, we conduct experiments on an extensive benchmark MME [21] that is tailored to evaluate MLLMs from multiple perspectives. As Fig. 5 shows, the fine-tuning with MMRel improves most MME tasks as compared with the baseline LLaVA-1.5 [39]. This experiment verifies the scalability and applicability of the proposed MMRel while conducting fine-tuning in various multi-modal tasks. Please refer to supplementary material for more details regarding the evaluation metrics and experimental results on the MME benchmark [21].
5 Conclusion
We notice that Multi-Modal Large Language Models (MLLMs) encounter difficulties in understanding inter-object relations, a research area that remains under-explored. To address this gap, we contribute the Multi-Modal Relation Understanding benchmark (MMRel) with a clear taxonomy of inter-object relations and the definition of relation hallucinations. MMRel is specifically designed to probe MLLMs’ capabilities on relation understanding and includes a challenging adversarial subset featuring highly unusual relations. Thanks to its large-scale, high-quality, and diverse data, MMRel serves as a versatile tool for both evaluating and fine-tuning MLLMs, enhancing their capabilities in relation understanding and other perceptual vision-language tasks. Extensive experiments and corresponding analysis validate the effectiveness of MMRel.